OpenStudy (seattle12345):

Can someone please help with this? Multiply out and write as a monomial. No negative exponents. (2xy^-3)^2(5y)^2

OpenStudy (anonymous):

Rules of exponents... 2^2 x^2 y^-6 * 5^2 y^2 = 4 x^2 y^-6 * 25 y^2 = 100 x^2 y^(-6+2) = 100 x^2 y^-4 = 100 x^2 / y^4

OpenStudy (seattle12345):

Thanks!